The fibonacci sequence is defined as x0 = 0, x1 = 1 and xn+2 = xn +xn+1 , for all non negative integers n prove that, xm = xr+1xm-r + xrxm-r for all integers m ≥ 1 and 0 ≤ r ≤ m-1 and xd divides xkd for all integers k and d
1
Expert's answer
2020-05-01T17:28:40-0400
  1. "x_m = x_{r+1}x_{m-r} + x_rx_{m-r}."

Let's simplify this statement:

"x_m=x_{r+1}x_{m-r} + x_rx_{m-r} = x_{m-r}(x_{r+1} + x_r) = x_{m-r}x_{r+2}" (*). Let's prove it using the induction by r and m simultaneously. Let's prove it for every m(so m is fixed).

Since "m\ge 1" , the base of induction is:

"m=1: x_1 = x_{1-r}x_{r+2}." Since "0 \le r \le m-1", we obtain that "r = 0" . Thus, we should check if"x_1 = x_{1-0}x_{0+2} = 1\cdot 1 = 1" . Hence, we proved the base of induction.

Now, we will assume, that the (*) is true for every "m<m_0". Let's prove it for "m= m_0":

Let's prove the base of induction(by r):

a) "r =0." Hence, we should check if: "x_{m_0} = x_{m_0 -0}x_{0+2} = x_{m_0}x_2 = x_{m_0}."

b) The induction assumption is that for every "r", such that: "0\le r < r_0 \le m_0 -1" , (*) is true. Let's prove it for "r = r_0." We should check the following:

"x_{m_0} = x_{m_0 - r_0}x_{r_0 +2} = x_{m_0 - r_0}x_{r_0 +1} + x_{m_0 - r_0}x_{r_0}." We can apply the induction assumption to both of the summands in the following way:

For the first summand:

"r_0 + 1 = (r_0 -1) + 2, m_0 -r_0 = (m_0 -1) - (r_0 -1)" , and we obtain the following(by using the (*)): "x_{m_0 - r_0}x_{r_0 +1} = x_{m_0 -1}". By applying the induction assumption and (*) to the second summand we obtain: "x_{m_0 - r_0}x_{r_0} = x_{m_0 -2}." Thus we should check if: "x_{m_0} = x_{m_0 -1} + x_{m_0 -2}." This is true.


2."x_d | x_{kd}," where "k,d" are some integers.

Let's use the previous statement:

"x_{kd} = x_{kd -d }x_{d+2} = x_{(k-1)d}x_{d+2}." By applying this fact "(k-1)" times, we obtain the following: "x_{kd} = x_{d} x_{d+2}^{k-1}."

Thus "(k \ge 1)" : "x_d |x_{kd}"